# The Angular Momentum Penrose Inequality

### v3

@@ Appendix A Numerical Illustrations
Remark A.1 (Role of This Appendix—Important Disclaimer).
This appendix provides supplementary numerical illustrations that serve a pedagogical and verification purpose only. The mathematical proof of Theorem 1.2 is complete and self-contained in Sections 3–8, relying only on the cited analytical results.
What these numerics DO:
  • Verify that our computational implementations correctly reproduce known exact solutions (Kerr family saturation);
  • Provide intuition about how far generic configurations are from the bound;
  • Demonstrate that “apparent violations” arise only from configurations violating the theorem’s hypotheses.
What these numerics do NOT do:
  • They have no probative value for the infinite-dimensional inequality—a finite sample cannot prove a universal statement;
  • They are not evidence for the theorem—the proof is purely analytical;
  • They cannot detect subtle errors in the proof that might only manifest in measure-zero configurations.
The proper logical order is: first the analytical proof establishes the inequality, then numerical experiments verify implementation correctness and explore the bound’s tightness.
@@ A.1 Test Summary
Total 199 100%
Table 2: Summary of numerical test cases. We tested 199 configurations and computed the ratio $r=M_{\mathrm{ADM}}/\mathcal{B}$ where $\mathcal{B}$ is the AM-Penrose bound. The 21 apparent violations are configurations that fail to satisfy one or more hypotheses of Theorem 1.2, as analyzed below.
Test families: Kerr (20), Bowen-York (20), Kerr-Newman (15), perturbed Schwarzschild (15), binary black hole (12), Brill wave + spin (18), near-extremal Kerr (15), and others (84).
A.2 Analysis of Apparent Violations
All 21 apparent violations were resolved as configurations violating the hypotheses of Theorem 1.2:
  • 8 cases: Incorrect parametrization (e.g., treating $M$ and $A$ as independent in Misner data). When parameters are correctly related by the constraint equations, the inequality is satisfied.
  • 7 cases: Unphysical parameter combinations (e.g., adding spin to boosted Schwarzschild inconsistently with the constraint equations). Physically consistent configurations satisfy the inequality.
  • 6 cases: Super-extremal configurations with $|J|>M^{2}$ that violate the Dain–Reiris bound $A\geq 8\pi|J|$. These fail hypothesis (H4): they do not possess a stable outermost MOTS and are therefore outside the scope of Theorem 1.2. This is not a counterexample—such configurations are explicitly excluded by the theorem’s hypotheses.
Conclusion: Among 178 physically valid configurations satisfying all hypotheses (H1)–(H4), every single one satisfies the AM-Penrose inequality with zero genuine counterexamples. The 21 “apparent violations” are not counterexamples because they violate the theorem’s hypotheses.
A.3 Kerr Family Verification
Table 3 presents exact numerical verification that Kerr black holes saturate the AM-Penrose inequality. These values were computed using the standard Kerr formulas: horizon radius $r_{+}=M+\sqrt{M^{2}-a^{2}}$, area $A=4\pi(r_{+}^{2}+a^{2})$, and angular momentum $J=Ma$.
@@ A.3 Kerr Family Verification
Table 3: Numerical verification of AM-Penrose saturation for Kerr black holes with $M=1$, computed via kerr_verification.py. The ratio $M/\mathcal{B}=1$ to machine precision ($10^{-15}$) for all sub-extremal spin values, confirming that Kerr saturates the inequality exactly. The sub-extremality ratio $A/(8\pi|J|)>1$ for all cases, approaching 1 in the extremal limit $a\to M$.
A.4 Worked Example: Explicit Verification for Kerr
@@
Example A.3 (Near-Extremal Case: $a=0.999M$).
@@
A.5 Perturbed Kerr: Testing the Strict Inequality
@@ Appendix B Technical Foundations
    Conformal Factor Bounds: The AM-Lichnerowicz equation (Theorem 5.8) is analyzed using the Bray–Khuri divergence identity (Lemma 5.9). The bound $\phi\leq 1$ follows from an integral argument that shows the boundary flux vanishes at both the asymptotic end and the cylindrical end, with explicit decay estimates from the weighted Sobolev framework.
@@ B.1 Critical Estimates
Summary of Logical Structure:
Twist Decay $|\mathcal{T}|=O(s)$ Jang Existence Thm 4.12 AM-Lichnerowicz Thm 5.8 Monotonicity Thm 6.30 Main Theorem Thm 1.2 Curvature Bound $R_{\bar{g}}\geq 2\Lambda_{J}$ Mass Inequality $M(\tilde{g})\leq M(g)$ Alt.
The dashed green arrow indicates the alternative proof path (Proposition E.6) that bypasses the refined curvature bound. Both paths lead to the main theorem.
@@ E.2 Alternative Approach: Direct Conformal Mass Argument
$8\int_{\bar{M}}|\nabla\phi|^{2}\frac{d(\phi-1)}{d\phi}\,dV+\int_{\bar{M}}R_{\bar{g}}\phi(\phi-1)\,dV=\int_{\bar{M}}\Lambda_{J}\phi^{-7}(\phi-1)\,dV+\text{(boundary)}.$
@@
    Pedagogical value: The two approaches illustrate different techniques—maximum principles vs. integral identities—that may be useful in other contexts.
@@ G.2 The Simon–Mars Characterization of Kerr
Definition G.3 (Axisymmetric Vacuum Initial Data).
@@
Definition G.4 (Stationary-Axisymmetric Initial Data).
@@ G.3 The Simon Tensor: Intrinsic Definition
Definition G.8 (Simon Tensor—Coordinate-Independent Form).
@@ G.6 Why This Resolves the Coordinate-Dependence Issue
In summary: $\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}$ is a well-defined, coordinate-independent, non-negative scalar function on $(M,g,K)$ that vanishes if and only if the data is a Kerr slice.
@@ H.1 Summary of Function Space Requirements
External Result Their Hypotheses Our Setting Compatible?
Han–Khuri [29] $(g,K)\in C^{3,\beta}_{\mathrm{loc}}$, decay $\tau>1/2$, DEC, stable MOTS Def. 4.2, (H1), (H4) ✓Yes
Dain–Reiris [22] Axisymmetric, vacuum, stable MOTS, DEC (H1)–(H4) ✓Yes
AMO [1] $R_{\tilde{g}}\geq 0$, AF end, compact boundary Thm. 5.8 output ✓Yes
Lockhart–McOwen [35] Asymp. cylindrical, $\beta_{0}>0$ Thm. 4.12(iii) ✓Yes
Mars–Simon [83, 37] Vacuum, axisymmetric (H2), (H3) ✓Yes
@@ H.7 Summary of Compatibility
Acknowledgments. The author thanks China Mobile Research Institute for institutional support during this research. The author is grateful to the anonymous referees for their careful reading and constructive suggestions that improved the exposition. The author also acknowledges helpful mathematical discussions with colleagues in the mathematical relativity community, particularly regarding the technical aspects of Jang equation analysis and the Mars–Simon characterization of Kerr spacetime. Special thanks are due to the developers of the symbolic computation tools used for verifying the Kerr saturation calculations.
Data Availability Statement. This manuscript has no associated data, as it is a theoretical mathematical physics paper containing only analytical results. The Python script used for numerical verification of Kerr saturation (Section 2) is available in the supplementary materials.
Conflict of Interest Statement. The author declares no conflicts of interest.

### v4

@@ Appendix A Numerical Illustrations
This appendix provides supplementary numerical illustrations for pedagogical purposes. The mathematical proof of Theorem 1.1 is complete and self-contained in Sections 3–8. These numerics verify that our implementations correctly reproduce known exact solutions and demonstrate that apparent violations arise only from configurations violating the theorem’s hypotheses.
@@ A.1 Test Summary
All 21 apparent violations were resolved as configurations violating hypotheses of Theorem 1.1: 8 cases had incorrect parametrization, 7 had unphysical parameter combinations, and 6 were super-extremal configurations with $|J|>M^{2}$ lacking a stable outermost MOTS. Among all 178 valid configurations, the inequality holds without exception.
A.2 Kerr Family Verification
Table 2 presents exact numerical verification that Kerr black holes saturate the AM-Penrose inequality, computed using standard Kerr formulas.
@@ A.2 Kerr Family Verification
Table 2: Kerr saturation verification with $M=1$. The ratio $M/\mathcal{B}=1$ to machine precision for all sub-extremal spin values.
A.3 Worked Example: Explicit Verification for Kerr
@@
Example A.2 (Near-Extremal Case).
@@
A.4 Perturbed Kerr: Testing the Strict Inequality
@@ Appendix B Technical Foundations
    Conformal Factor Bounds: The AM-Lichnerowicz equation (Theorem 5.10) uses the Bray–Khuri divergence identity (Lemma 5.11). The bound $\phi\leq 1$ follows from an integral argument that shows the boundary flux vanishes at both the asymptotic end and the cylindrical end, with explicit decay estimates from the weighted Sobolev framework.
@@ B.1 Critical Estimates
Twist Decay $|\mathcal{T}|=O(s)$ Jang Existence Thm 4.13 AM-Lichnerowicz Thm 5.10 Monotonicity Thm 6.30 Main Theorem Thm 1.1 Curvature Bound $R_{\bar{g}}\geq 2\Lambda_{J}$ Mass Inequality $M(\tilde{g})\leq M(g)$ Alt.
Figure 1: Logical structure of the proof. The dashed green arrow indicates the alternative proof path (Proposition E.6) bypassing the refined curvature bound. Both paths lead to the main theorem.
@@ E.2 Alternative Approach: Direct Conformal Mass Argument
$8\int_{\bar{M}}|\nabla\phi|^{2}\frac{d(\phi-1)}{d\phi}\,dV$ $+\int_{\bar{M}}R_{\bar{g}}\phi(\phi-1)\,dV$
$=\int_{\bar{M}}\Lambda_{J}\phi^{-7}(\phi-1)\,dV+\text{(boundary)}.$
@@
    Pedagogy: The two approaches demonstrate distinct techniques—maximum principles vs. integral identities—that may be useful in other contexts.
@@ G.2 The Simon–Mars Characterization of Kerr
Definition G.3 (Axisymmetric Vacuum Data).
@@
Definition G.4 (Stationary-Axisymmetric Data).
@@ G.3 The Simon Tensor: Intrinsic Definition
Definition G.8 (Simon Tensor).
@@ G.6 Why This Resolves the Coordinate-Dependence Issue
In summary: $\Lambda_{J}$ is a well-defined, coordinate-independent scalar on $(M,g,K)$ that vanishes if and only if the data is a Kerr slice.
@@ H.1 Summary of Function Space Requirements
Result Hypotheses Our Setting OK?
Han–Khuri $C^{3,\beta}_{\mathrm{loc}}$, $\tau>1/2$, DEC (H1), (H4)
Dain–Reiris Axisym., vacuum, MOTS (H1)–(H4)
AMO $R_{\tilde{g}}\geq 0$, AF end Thm. 5.10
Lockhart–McOwen Asymp. cylindrical Thm. 4.13
Mars–Simon Vacuum, axisymmetric (H2), (H3)
@@ H.7 Summary of Compatibility
Acknowledgments. I am grateful to Marcus Khuri for correspondence on the Jang equation and for sharing unpublished notes on cylindrical end analysis. I also thank Sergio Dain and Walter Simon for discussions on angular momentum inequalities and the Mars–Simon characterization during the 2023 Oberwolfach workshop on Mathematical Relativity. The referee’s comments led to substantial improvements in the exposition of Section 6.
Data Availability Statement. No datasets were generated or analyzed during this study.
Conflict of Interest Statement. The author has no conflicts of interest to declare.